OpenStudy (anonymous):

the cruising speed of the bullet train will be no less than 170 mph. Use s to represent the trains cruising speed (in mph)

OpenStudy (anonymous):

\[s \ge 170\]

OpenStudy (anonymous):

the distance to the nearest exit door is less than 150 ft. Use d to represent the distance (in feet) to the nearest exit

OpenStudy (anonymous):

\[d < 150\]

OpenStudy (anonymous):

to be eligible for financial aid a student must earn no more than $25,000 per year. Use S to represent the yearly earnings (in dollars) of a student eligible for financial aid @koook

OpenStudy (anonymous):

to qualify for the championship, a runner must complete the race in at most 40 minutes. Use T to represent the time (in minutes) of a runner who qualifies for chmpionship

OpenStudy (anonymous):

\[s \le $25000\]

OpenStudy (anonymous):

thanks for helping I appreciate it

OpenStudy (anonymous):

\[t \le 40\]

OpenStudy (anonymous):

no problem. medal?

OpenStudy (anonymous):

I have more but sure

OpenStudy (anonymous):

the vehicles fuel efficiency is no less than 35 miles per gallon. Use F to represent the fuels efficiency (in miles per gallon)

OpenStudy (anonymous):

at most \[is \le \] at least \[is \ge \] more than or greater than \[is > \] less than or fewer than \[is < \]

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

\[F \ge 35\]